OpenStudy (anonymous):

Write these series with summation notation. 1) 1+4+7+ . . . +19 2) 1, 4, 9, 16, . . .

OpenStudy (anonymous):

\[\sum_{i=1}^{19} i+3 \]

OpenStudy (anonymous):

2) is not a series.

OpenStudy (amistre64):

its a seq of perfect squares tho

OpenStudy (anonymous):

For 2)?

OpenStudy (anonymous):

Yes, we all can see \(8^2\) in amistre64

OpenStudy (amistre64):

woohoo!! i thought i was the only one :)

OpenStudy (anonymous):

lol, you are taken after all.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

mugwump101 answer is incorrect.

OpenStudy (anonymous):

I think you can write number 2 as a series though, can't you?

OpenStudy (anonymous):

Really? why?

OpenStudy (amistre64):

you cant sum a series that isnt a series

OpenStudy (amistre64):

there aint 19 terms

OpenStudy (amistre64):

1 to ..... not 19

OpenStudy (amistre64):

16+3 = 19; 19+3 not= 19

OpenStudy (amistre64):

and 2+3 = 5 not 4

OpenStudy (amistre64):

its just all messed up lol

OpenStudy (anonymous):

Here is the answer \( \large \sum \limits_{n=1}^{7} 3n-2 \)

OpenStudy (amistre64):

\[\sum_{i=1}^{}-2+3(i)\]yeah that thinkg

OpenStudy (anonymous):

That makes alot more sense, thanks fellas. :)

OpenStudy (anonymous):

amstre you are slow :P

OpenStudy (amistre64):

its this accursed IE im playing n; and my old brain i have to work with too lol

OpenStudy (anonymous):

LOL I see. Thanks for your input XD